OpenStudy (anonymous):

Help.. Solve the equation by extracting square root. (x-7)2=(x+3)2

OpenStudy (anonymous):

x= 2

OpenStudy (anonymous):

Yes, that's what the book has, but how did you get it? Thanks for the reply.

OpenStudy (anonymous):

u welcome.... to understand it, u have to talk to me on skype....... nothing to pay.. r u ready.......

OpenStudy (anonymous):

to be honest i've never skyped before.

OpenStudy (anonymous):

when you square root a number you can get a positive and negative number. so we will use the negative number because when you use the positive the x eliminate it self. so here the solution: \[x - 7 = -(x+3)\] \[x-7= -x-3\] \[2x = 7 - 3\] \[2x = 4\]

OpenStudy (anonymous):

x= 2

OpenStudy (anonymous):

tinamrod, u can use it, it is easy..........

OpenStudy (anonymous):

@corec. thanks. I will try to work it out. @sudhamshu. someone just sent me the solution. Thanks again.!!

OpenStudy (anonymous):

Your welcome!

OpenStudy (anonymous):

tinamrod........ i just want to ask a question......... what do u want? Do u want to know the way a particular person solve your probelm? or Do u want to be able that u can solve such type of problem my yourself ? think .............

OpenStudy (anonymous):

I will work the problem out. If I need any help. I will look you up.

OpenStudy (anonymous):

this question is not only for u........., it is an important question for those who want to learn math??????????? my math teacher said second question is more important than first..!!!!!